A Carpenter is building a rectangular room with a fixed perimeter of 192ft. What dimensions would yield the maximum area? What is the maximum area?
the length that would yield the maximum area is ________ ft.

Answer by stanbon(75887)

A Carpenter is building a rectangular room with a fixed perimeter of 192ft.
Perimeter = 2(L + W)
192 = 2(L+W)
L + W = 96
----
What dimensions would yield the maximum area? What is the maximum area?
the length that would yield the maximum area is ________ ft.
---
Area = LW
A = L(96-L)
A = -L^2+96L
---
Quadratic with a = -1 and b = 96
-----
Maximum Area occurs when L = -b/(2a) = -96/(2*-1) = 48
----
If L = 48 ft. and L+W = 96, W = 48 ft.
-------
Maximum Area = 48*48 = 2304 sq. ft.
=====================================
Cheers,
Stan H.
===============